OpenStudy (anonymous):

how would I sketch the curve consisting of the points (x,y) such that e^(4y^2)e^(4y)e^(-x)=e^2

OpenStudy (dumbcow):

you can rewrite this as: \[e^{4y^2 +4y-x} = e^2\] then take log of both sides \[4y^2 +4y-x = 2\] isolate x \[x = 4y^2 +4y-2\] this is a sideways parabola complete the square to get in vertex form then graph

OpenStudy (dumbcow):

\[x = 4(y+\frac{1}{2})^2 -3\] vertex at (-3, -1/2)

OpenStudy (larseighner):

\[\large e^{4y^2}e^{4y}e^{-x}=e^2\] \[\large e^{4y^2}e^{4y}={e^2 \over e^{-x}} \] \[\large e^{4y^2}e^{4y}=e^2 e^x \] Now I argue from the law of exponents, which works with any base, that this implies the equation that @dumbcow found, and the solution proceeds a @dumbcow wrote.